# Fluctuations and the limit of predictability in protein evolution

### v2

@@ III.1 Mutational dynamics and its fluctuations
III.2 Dynamical heterogeneity across families
@@
For compactness, we only display the average dynamical fluctuations, with $\chi_{\rm tot}$ corresponding to the total variance of $H$, $\chi_{\rm dyn}$ to the variance due to the stochatisticy of the evolution (averaged over the ancestor), and $\chi_{\rm bg}$ to the variance due to the choice of ancestor. All the $\chi_{\rm tot}$ curves display a similar behavior, increasing from zero to a maximum value that is maintained over one or more decades, before finally decreasing to a smaller value corresponding to equilibrium. What matters most to us, however, is the relative importance of the two terms $\chi_{\rm dyn}$ and $\chi_{\rm bg}$ in which $\chi_{\rm tot}$ can be decomposed. The key observation is that background-related fluctuations dominate over dynamical ones at short evolutionary times, while the inverse is true at larger times. This means that, at least over time scales for which $\chi_{\rm bg}\gg\chi_{\rm dyn}$, we can hope to reconstruct with good accuracy the ancestral sequence from which a set of evolutionary trajectories started. Conversely, at larger time scales, the dynamical noise contribution dominates and the trajectory-to-trajectory fluctuations are large enough to hide the signal coming from the ancestral sequence, precluding the possibility to reconstruct it. This behavior is observed in all the protein families that we tested, with the exception of the WW domain in which $\chi_{\rm bg}$ never grows to larger values than $\chi_{\rm dyn}$. We attribute this difference to the small size of the WW domain, which prevents the sequences to accumulate enough epistatic interactions. However, both the time scales and the relative importance of the two terms contributing to the total susceptibility vary greatly from family to family (Fig. 3). In DBD, AAC6, and DHFR the peak of $\chi_{\rm tot}$ is reached quite early in the dynamics and the two contributions $\chi_{\rm dyn}$ and $\chi_{\rm bg}$ are almost non-overlapping. On the other hand, for CM, BL, and SP (the longer families we analyzed) the peak occurs much later due to the two contributions having a significant overlap. Overall, these results point to an increasing complexity of the fitness landscape with the length of the sequence, as intuitively expected.
@@
III.3 Epistatic constraints and dynamical fluctuations
@@
In Fig. 4 we show, for the three ancestral sequences highlighted in Fig. 2, how the dynamical part of the susceptibility is strongly related to the epistatically constrained sites (defined as in Sec. II.2 and in Ref. Bari et al. (2024)). To make this relation clear, we first checked that the time scale at which the dynamical susceptibility reaches its peak is compatible with the time scale at which epistatically constrained sites evolve. For a given ancestor $A_{0}$, we consider the five most epistatically constrained sites, i.e. those with the largest ${\rm CIE}_{i}-{\rm CDE}_{i}^{A_{0}}$. For these sites, we consider at time $t$ the frequency of appearance of amino acid $a$, $f_{i}^{t}(a)$, in the set of sequences that evolved from $A_{0}$, and from it we compute a time-dependent entropy
@@
To quantify these observations, in Fig. 5 we consider the same set of families as in Fig. 3, namely WW, DBD, CM, AAC6, DHFR, BL, and SP, spanning a wide range of sequence length (from $L=31$ in WW to $L=254$ in SP). For each family, we consider a set of ancestors $A_{0}$ and we report a scatter plot of the following quantities, each $A_{0}$ being a point:
@@
III.4 Cooperative mutational dynamics
@@
where $\langle\dots\rangle_{\rm nat}$ is the average computed over the natural sequences. We checked that this is indeed the case.
In section III.3 we showed that the presence of a peak in $\chi^{A_{0}}_{\rm dyn}$ is correlated to $\textrm{CDE}^{A_{0}}$, i.e. to the amount of epistatically constrained sites in the ancestral sequence. In this section, we show that these sites also display large dynamical correlations between themselves. The bar above each plot in Fig. 6 is shaded to represent $\textrm{CIE}_{i}-\textrm{CDE}^{\rm A}_{i}$, capturing the information about site $i$ encoded in the background sequence $A_{\setminus i}$. This coloring reflects the degree to which each site is constrained by epistatic interactions. The conserved and variable sites are highlighted by a bright yellow, while the epistatically constrained ones are indicated by a dark blue. As expected, we see good agreement between these residues and the ones that give a large contribution to $\chi^{A_{0}}_{\rm dyn}$. Hence, we conclude that epistatically constrained sites mutate cooperatively around the time scale corresponding to the peak of $\chi^{A_{0}}_{\rm dyn}$.
III.5 Response to environmental variation
In the literature on disordered physical systems, it has been established that the dynamical correlations discussed in sections III.3 and III.4 are related to the linear response of the dynamics to a change in temperature through a kind of fluctuation-dissipation relation Berthier et al. (2005).
@@ III.5 Response to environmental variation
The results in Fig. 7 suggest that more epistatically constrained ancestral sequences, which display a stronger peak in the dynamical susceptibility, will also display a stronger response of the evolutionary dynamics to a small change in selection strength. Such response is stronger around the time of the peak in $\chi^{A_{0}}_{\rm dyn}(t)$, which is also the time at which epistatically constrained sites mutate cooperatively, see section III.4. In other words, sequences with more epistatically constrained sites are more sensitive to environmental changes. In the SM, we give an additional analytical argument based on the short-time dynamics, that supports the same conclusion.
III.6 Limits of ancestral sequence reconstruction
In an evolutionary process, predictions can be made in two directions: either starting from the ancestral sequence and predicting its future evolution over a given time; or starting from a set of evolved sequences and attempting to infer the ancestral sequence from which the evolution began. An attempt at predicting future evolution has been made in Ref. Bisardi et al. (2022) (see Widmer-Cooper et al. (2008); Schoenholz et al. (2016); Bapst et al. (2020); Jung et al. (2023, 2024) for related studies in disordered systems), where however the prediction was limited to small evolutionary time scales. Because we have repeatedly hinted that our analysis can provide insight into the time scale over which reconstructing the ancestral sequence of an evolutionary process is possible, before concluding, we consider here the ancestral sequence reconstruction problem more explicitly.
@@ III.6 Limits of ancestral sequence reconstruction
The results are shown in Fig. 8a, where we plot the evolution of the average Hamming distance for the same sequences as in Fig. 2, and we compare it with the Hamming distance (i.e. the percentage of errors) between the ground truth ancestor and the reconstructed one. At the beginning of the dynamics the FastML tool is able to reconstruct perfectly the original sequence in all of the three cases. This is not surprising, as at short times few mutations appear. The situation is similar at large times, where the MSAs generated from the three starting sequences are statistically indistinguishable and the FastML tool performs equally bad in all of them. Thus, it is more insightful to look at intermediate times, where a significant difference between the evolving sequences is present. After $\sim 10$ sweeps the variable sites begin to mutate and the ancestral sequence reconstruction tool cannot reproduce the initial sequence perfectly. The leap in the error percentage with respect to the previous point is however larger for the blue sequence, which is the one with fewer epistastic sites, while the red one is still recovered reasonably well. This difference is present for the subsequent points as well, until the error saturates. In Fig. 8b we show that the Hamming distance (percentage of errors) of the sequence reconstruction results grows roughly proportionally to the average sequence divergence of the evolutionary trajectories for all of the three studied sequences, but more so for the more epistatic red sequence. We conclude that, at least with the FastML tool, the possibility of reconstructing ancestral sequences is mostly a matter of divergence, with more errors accumulating when using datasets of more diverged sequences.
@@
Our analysis shows that the amount of diversity at a given evolutionary time depends strongly on the ancestor, due to highly non-trivial epistatic dynamical correlations. More epistatically constrained ancestors give rise to less diversity, thus allowing for reconstruction over longer evolutionary times (Fig. 8a). Yet, at comparable amount of diversity, more epistatic ancestors are more difficult to reconstruct (Fig. 8b), at least using the FastML algorithm that neglects correlations between sites Felsenstein (2003). These results confirm once more that the evolutionary time limit for predictability is linked to the presence of epistatic sites that carry information about the starting sequence only up to a certain sequence divergence. We further observe that more advanced tools (yet to be fully developed) for ancestral sequence reconstruction might exploit the epistatic correlations to achieve a better reconstruction De Leonardis et al. (2024), which is not done in FastML and similar algorithms. We expect that such tools would perform better on more epistatic ancestors such as the red one in Fig. 8.
IV Conclusions
@@
In this study, we explore the role of disorder and stochasticity in protein evolution, focusing on the interplay between epistasis, site correlations, and sequence-dependent fluctuations. We simulated evolution in silico using DCA to model the fitness landscape, and a MCMC dynamics to mimic the process of mutation and selection, a methodology that has been validated in previous work de la Paz et al. (2020); Bisardi et al. (2022); Bari et al. (2024). By employing tools from statistical physics Kirkpatrick and Thirumalai (1988); Franz et al. (1999); Franz and Parisi (2000); Bouchaud and Biroli (2005); Berthier et al. (2005); Franz et al. (2011); Berthier et al. (2011); Franz et al. (2013); Seoane and Zamponi (2018); Berthier and Jack (2007), we quantified how the ancestral amino acid sequence significantly influences early evolutionary dynamics, as measured by the dynamical susceptibility. Our main results are the following.
@@
  • Our analysis shows that, during the initial phase of evolution, there is very significant heterogeneity in the dynamics, depending on the choice of ancestral sequence. Some ancestral sequences lead to a smooth evolution where each site mutates almost independently of the others, while others lead to a more complex dynamics characterized by intermediate plateaus and significant fluctuations (Sec. III.1).
@@
  • We have shown that, for a variety of protein families, the noise arising from the starting sequence dominates over stochastic evolutionary fluctuations, implying that one can, in principle, trace the evolutionary trajectory back to its origin (Sec. III.2). This effect is more or less pronounced, depending on the family, and our tools allow one to quantify it. However, as time progresses and epistatically constrained sites evolve, this traceability is lost due to the growing influence of the mutational stochasticity.
@@
  • The heterogeneity of the initial evolutionary dynamics can be traced back to the amount of epistatic constraints in the ancestral sequence. We introduced a quantity, the average over sites of the context-dependent entropy, and we have shown that this quantity is strongly correlated with the time at which dynamical heterogeneity reaches its peak and with the strength of the fluctuations at the peak (Sec. III.3).
@@
  • The amplitude of the global fluctuations can be expressed in terms of a sum of pairwise dynamical correlations between pairs of residues. More epistatically constrained ancestral sequences show groups of residues that mutate collectively, and those patterns can be identified from the correlation matrix $G_{ij}^{A_{0}}(t)$ we introduced in Sec. III.4. The observed correlations between sites thus reflect a complex epistatic landscape, where certain residues evolve in a highly context-dependent manner.
@@
  • We demonstrated that these epistatic sites, which evolve over long time scales, significantly affect the response of the dynamics to a change in environmental conditions. More epistatically constrained sequences lead to a larger response to a change in environment, suggesting potential implications on sequence evolvability imposed by environmental pressures, such as antibiotic concentration (Sec. III.5).
@@
More generally, our findings extend the analogy between protein evolution and disordered physical systems, reinforcing the idea that protein dynamics exhibit characteristics of complex and strongly correlated systems. However, when comparing these results with statistical physics models exhibiting glassy dynamics, we also found qualitative differences, emphasizing the unique constraints imposed by natural evolution on proteins. Future work should aim to further elucidate the connection between epistasis, evolvability, and environmental selection, which may offer insights into evolutionary strategies across diverse biological systems. Furthermore, we believe that most of the ideas presented in this work will soon be amenable to experimental testing, thanks to the increased power of in vitro evolution platforms.
@@ Acknowledgements.
  • SP: the alignment comes from a private communication with the group of Clément Nizak in Paris.
@@
As we already pointed out, the evolutionary protocol that we implemented is a simple Metropolis Monte Carlo defined over amino acid sequences, which is far from being realistic. A more detailed approach was recently described in Bari et al. (2024), where it was shown that an algorithm working with single-nucleotide substitutions, insertions and deletions was able to correctly reproduce the dynamics both at long and short time scales. We checked that this more refined algorithm, which was shown to over-perform a simple Metropolis at short time scales Bisardi et al. (2022), produces qualitatively similar results for the quantities of interest for our work (Fig. S1b).

### v3

@@ IV.1 Mutational dynamics and its fluctuations
IV.2 Dynamical heterogeneity across families
@@
For compactness, we only display the average dynamical fluctuations, with $\chi_{\rm tot}$ corresponding to the total variance of $H$, $\chi_{\rm dyn}$ to the variance due to the stochatisticy of the evolution (averaged over the ancestor), and $\chi_{\rm bg}$ to the variance due to the choice of ancestor. All the $\chi_{\rm tot}$ curves display a similar behavior, increasing from zero to a maximum value that is maintained over one or more decades, before finally decreasing to a smaller value corresponding to equilibrium. What matters most to us, however, is the relative importance of the two terms $\chi_{\rm dyn}$ and $\chi_{\rm bg}$ in which $\chi_{\rm tot}$ can be decomposed. The key observation is that background-related fluctuations dominate over dynamical ones at short evolutionary times, while the inverse is true at larger times. This means that, at least over time scales for which $\chi_{\rm bg}\gg\chi_{\rm dyn}$, we can hope to reconstruct with good accuracy the ancestral sequence from which a set of evolutionary trajectories started. Conversely, at larger time scales, the dynamical noise contribution dominates and the trajectory-to-trajectory fluctuations are large enough to hide the signal coming from the ancestral sequence, precluding the possibility to reconstruct it. This behavior is observed in all the protein families that we tested, with the exception of the WW domain in which $\chi_{\rm bg}$ never grows to larger values than $\chi_{\rm dyn}$. We attribute this difference to the small size of the WW domain, which prevents the sequences to accumulate enough epistatic interactions. However, both the time scales and the relative importance of the two terms contributing to the total susceptibility vary greatly from family to family (Fig. 3). In DBD, AAC6, DHFR, and SP the peak of $\chi_{\rm tot}$ is reached quite early in the dynamics and the two contributions $\chi_{\rm dyn}$ and $\chi_{\rm bg}$ are almost non-overlapping. On the other hand, for CM and BL, the peak occurs much later due to the two contributions having a significant overlap.
@@
IV.3 Epistatic constraints and dynamical fluctuations
@@
In Fig. 4 we show, for the three ancestral sequences highlighted in Fig. 2, how the dynamical part of the susceptibility is strongly related to the epistatically constrained sites (defined as in Sec. III.2 and in Ref. Bari et al. (2024)). To make this relation clear, we first checked that the time scale at which the dynamical susceptibility reaches its peak is compatible with the time scale at which epistatically constrained sites evolve. For a given ancestor $A_{0}$, we consider the five most epistatically constrained sites, i.e. those with the largest ${\rm CIE}_{i}-{\rm CDE}_{i}^{A_{0}}$. For these sites, we consider at time $t$ the frequency of appearance of amino acid $a$, $f_{i}^{t}(a)$, in the set of sequences that evolved from $A_{0}$, and from it we compute a time-dependent entropy
@@
To quantify these observations, in Fig. 5 we consider the same set of families as in Fig. 3, namely WW, DBD, CM, AAC6, DHFR, BL, and SP, spanning a wide range of sequence length (from $L=31$ in WW to $L=220$ in SP). For each family, we consider a set of ancestors $A_{0}$ and we report a scatter plot of the following quantities, each $A_{0}$ being a point:
@@
IV.4 Cooperative mutational dynamics
@@
where $\langle\dots\rangle_{\rm nat}$ is the average computed over the natural sequences.
In section IV.3, we showed that the presence of a peak in $\chi^{A_{0}}_{\rm dyn}$ is correlated to $\textrm{CDE}^{A_{0}}$, i.e. to the amount of epistatically constrained sites in the ancestral sequence. In this section, we show that these sites also display large dynamical correlations between themselves. The bar above each plot in Fig. 6 is shaded to represent $\textrm{CIE}_{i}-\textrm{CDE}^{\rm A}_{i}$, capturing the information about site $i$ encoded in the background sequence $A_{\setminus i}$. This coloring reflects the degree to which each site is constrained by epistatic interactions. The conserved and variable sites are indicated in white, while the epistatically constrained ones are shown in black. As expected, we see good agreement between these residues and the ones that give a large contribution to $\chi^{A_{0}}_{\rm dyn}$. Hence, we conclude that epistatically constrained sites mutate cooperatively around the time scale corresponding to the peak of $\chi^{A_{0}}_{\rm dyn}$.
IV.5 Response to environmental variation
In the literature on disordered physical systems, it has been established that the dynamical correlations discussed in sections IV.3 and IV.4 are related to the linear response of the dynamics to a change in temperature through a kind of fluctuation-dissipation relation Berthier et al. (2005).
@@ IV.5 Response to environmental variation
The results in Fig. 7 suggest that more epistatically constrained ancestral sequences, which display a stronger peak in the dynamical susceptibility, will also display a stronger response of the evolutionary dynamics to a small change in selection strength. Such response is stronger around the time of the peak in $\chi^{A_{0}}_{\rm dyn}(t)$, which is also the time at which epistatically constrained sites mutate cooperatively, see section IV.4. In other words, sequences with more epistatically constrained sites are more sensitive to environmental changes. In the SM, we give an additional analytical argument based on the short-time dynamics, that supports the same conclusion.
IV.6 Limits of ancestral sequence reconstruction
In an evolutionary process, predictions can be made in two directions: either starting from the ancestral sequence and predicting its future evolution over a given time; or starting from a set of evolved sequences and attempting to infer the ancestral sequence from which the evolution began. An attempt at predicting future evolution has been made in Ref. Bisardi et al. (2022) (see Refs. Widmer-Cooper et al. (2008); Schoenholz et al. (2016); Bapst et al. (2020); Jung et al. (2025, 2024) for related studies in disordered systems), where however the prediction was limited to small evolutionary time scales. Because we have repeatedly hinted that our analysis can provide insight into the time scale over which reconstructing the ancestral sequence of an evolutionary process is possible, before concluding, we consider here the ancestral sequence reconstruction problem more explicitly.
@@ IV.6 Limits of ancestral sequence reconstruction
The results are shown in Fig. 8a, where we plot the evolution of the average Hamming distance for the same sequences as in Fig. 2, and we compare it with the Hamming distance (i.e. the percentage of errors) between the ground truth ancestor and the reconstructed one. At the beginning of the dynamics the FastML tool is able to reconstruct perfectly the original sequence in all of the three cases. This is not surprising, as at short times few mutations appear. The situation is similar at large times, where the MSAs generated from the three starting sequences are statistically indistinguishable and the FastML tool performs equally bad in all of them Gascuel and Steel (2010); Evans et al. (2000). Thus, it is more insightful to look at intermediate times, where a significant difference between the evolving sequences is present. After $\sim 10$ sweeps the variable sites begin to mutate and the ancestral sequence reconstruction tool cannot reproduce the initial sequence perfectly. The leap in the error percentage with respect to the previous point is however larger for the blue sequence, which is the one with fewer epistastic sites, while the red one is still recovered reasonably well. This difference is present for the subsequent points as well, until the error saturates. In Fig. 8b we show that the Hamming distance (percentage of errors) of the sequence reconstruction results grows roughly proportionally to the average sequence divergence of the evolutionary trajectories for all of the three studied sequences, but more so for the more epistatic red sequence. We conclude that, at least with the FastML tool, the possibility of reconstructing ancestral sequences is mostly determined by the Hamming distance of the ancestor from the evolving sequences, i.e. $[H^{A_{0}}(t)]$, with more errors accumulating when using datasets of more diverged sequences.
@@
Our analysis shows that, for a given evolutionary time, the amount of diversity $[H^{A_{0}}(t)]$ depends strongly on the ancestor, due to highly non-trivial epistatic dynamical correlations. More epistatically constrained ancestors give rise to less diversity, thus allowing for reconstruction over longer evolutionary times (Fig. 8a). Yet, at comparable amount of diversity, more epistatic ancestors are more difficult to reconstruct (Fig. 8b), at least using the FastML algorithm that neglects correlations between sites Felsenstein (2003). These results confirm once more that the evolutionary time limit for predictability is linked to the presence of epistatic sites that carry information about the starting sequence only up to a certain sequence divergence. We further observe that more advanced tools (yet to be fully developed) for ancestral sequence reconstruction might exploit the epistatic correlations to achieve a better reconstruction De Leonardis et al. (2025), which is not done in FastML and similar algorithms. We expect that such tools would perform better on more epistatic ancestors such as the red one in Fig. 8.
V Conclusions
@@
In this study, we explore the role of disorder and stochasticity in protein evolution, focusing on the interplay between epistasis, site correlations, and sequence-dependent fluctuations. We simulated evolution in silico using DCA to model the fitness landscape, and an MCMC algorithm to mimic the process of mutation and selection, a methodology that has been validated in previous work de la Paz et al. (2020); Bisardi et al. (2022); Bari et al. (2024). By employing tools from statistical physics Kirkpatrick and Thirumalai (1988); Franz et al. (1999); Franz and Parisi (2000); Bouchaud and Biroli (2005); Berthier et al. (2005); Franz et al. (2011); Berthier et al. (2011); Franz et al. (2013); Seoane and Zamponi (2018); Berthier and Jack (2007), we quantified how the ancestral amino acid sequence significantly influences early evolutionary dynamics, as measured by the dynamical susceptibility. Our main results are the following.
@@
  • Our analysis shows that, during the initial phase of evolution, there is very significant heterogeneity in the dynamics, depending on the choice of ancestral sequence. Some ancestral sequences lead to a smooth evolution where each site mutates almost independently of the others, while others lead to a more complex dynamics characterized by intermediate plateaus and significant fluctuations (Sec. IV.1).
@@
  • We have shown that, for a variety of protein families, the noise arising from the starting sequence dominates over stochastic evolutionary fluctuations, implying that one can, in principle, trace the evolutionary trajectory back to its origin (Sec. IV.2). This effect is more or less pronounced, depending on the family, and our tools allow one to quantify it. However, as time progresses and epistatically constrained sites evolve, this traceability is lost due to the growing influence of the mutational stochasticity.
@@
  • The heterogeneity of the initial evolutionary dynamics can be traced back to the amount of epistatic constraints in the ancestral sequence. We introduced a quantity, the average over sites of the context-dependent entropy, and we have shown that this quantity is strongly correlated with the time at which dynamical heterogeneity reaches its peak and with the strength of the fluctuations at the peak (Sec. IV.3). This method for assessing sequence evolvability could guide experimentalists in selecting an appropriate starting sequence for experiments, based on the desired functional behavior.
@@
  • The amplitude of the global fluctuations can be expressed in terms of a sum of pairwise dynamical correlations between pairs of residues. More epistatically constrained ancestral sequences show groups of residues that mutate collectively, and those patterns can be identified from the correlation matrix $G_{ij}^{A_{0}}(t)$ we introduced in Sec. IV.4. The observed correlations between sites thus reflect a complex epistatic landscape, where certain residues evolve in a highly context-dependent manner. These correlations could also be leveraged in experimental settings; for instance, one may imagine to control the evolvability of a specific residue by targeting sites that exhibit strong correlations with it.
@@
  • We demonstrated that these epistatic sites, which evolve over long time scales, significantly affect the response of the dynamics to a change in environmental conditions. More epistatically constrained sequences lead to a larger response to a change in environment, suggesting potential implications on sequence evolvability imposed by environmental pressures, such as antibiotic concentration (Sec. IV.5).
@@
More generally, our findings extend the analogy between protein evolution and disordered physical systems, reinforcing the idea that protein dynamics exhibit characteristics of complex and strongly correlated systems. However, when comparing these results with statistical physics models exhibiting glassy dynamics, we also found qualitative differences, emphasizing the unique constraints imposed by natural evolution on proteins. Future work should aim to further elucidate the connection between epistasis, evolvability, and environmental selection, which may offer insights into evolutionary strategies across diverse biological systems. Moreover, more realistic evolutionary dynamics could be considered. In this paper we focused on independent Monte Carlo chains, which corresponds to evolution on a star tree. Introducing a more complex tree structure could affect the results presented here, as the varying distance between the leaves of the tree and its root and the correlations coming from a common ancestor may affect the computation of $\chi^{A_{0}}_{\rm dyn}$. Furthermore, we believe that most of the ideas presented in this work will soon be amenable to experimental testing, thanks to the increased power of in vitro evolution platforms.
@@ Acknowledgements.
  • SP: we downloaded the alignment corresponding to the PF00089 family from Pfam Mistry et al. (2020) and excluded sequences with more than 20% gaps.
@@
As we already pointed out, the evolutionary protocol that we implemented is a simple Metropolis Monte Carlo defined over amino acid sequences, which is schematically described as follows:
  1. 1.
    Initialize with a protein sequence $A^{0}$.
  2. 2.
    At each iteration:
    1. (a)
      Select a random site $i\in\{1,\dots,L\}$.
    2. (b)
      Propose a mutation $a_{i}^{\prime}\neq a_{i}$ by selecting a new amino acid at site $i$ uniformly from the alphabet.
    3. (c)
      Compute the energy difference:
      $\Delta H=H(A^{\prime})-H(A)\ ,$ (S1)
      where $A^{\prime}$ is the sequence with the proposed mutation.
    4. (d)
      Accept the mutation with probability
      $P_{\text{accept}}=\min(1,e^{-\Delta H/T})\ .$ (S2)
      If accepted, update $a_{i}\leftarrow a_{i}^{\prime}$; otherwise, keep $a_{i}$.
  3. 3.
    Repeat for a sufficiently large number of iterations to reach equilibrium.
After equilibration, sequences sampled from this procedure approximate the equilibrium distribution of the Potts model.
This dynamics is far from being realistic, especially at short time-scales. As a matter of fact, a more detailed approach was recently described in Bari et al. (2024), where it was shown that an algorithm working with single-nucleotide substitutions (instead of amino acid substitutions), insertions and deletions was able to correctly reproduce the dynamics both at long and short time scales. We checked that this more refined algorithm, which was shown to over-perform a simple Metropolis at short time scales Bisardi et al. (2022), produces qualitatively similar results for the quantities of interest for our work (Fig. S1b).